OpenStudy (anonymous):

The current student population of the Brentwood Student Center is 2,000. The enrollment at the center increases at a rate of 4% each year. To the nearest whole number, what will the student population be closest to in 3 years? (1) 2,240 (3) 5,488 (2) 2,250 (4) 6,240

OpenStudy (anonymous):

i don't understand these type of question im lost!!!

OpenStudy (anonymous):

Find population growth in first year. 2000 x 4% = 80 Add this to initial population. 2000 + 80 = 2080 = population in 1 year Find population growth in second year. 2080 x 4% = 83.2 Add this to year 2 population. 2080 + 83.2 = 2163.2 = population in 2 years Find population growth in third year. 2163.2 x 4% = 86.5 (approximate) Add this to year 3 population. 2163.2 + 86.5 = 2249.7 = population in 3 years Round the answer to 2,250.

OpenStudy (anonymous):

thank you so much!! :)

OpenStudy (anonymous):

wow you can also try this to increase a number by \(4\%\) multiply it by \(1.04\) to increase it by \(4\%\) three times in a row, multiply it by \((1.04)^3\)

OpenStudy (anonymous):

gonna save you a lot of work if you have to find the population in ten years rather than in 3